OpenStudy (mkmkasim):

what is the antiderivative ∫(5^x)(sin(5^x))dx

OpenStudy (mkmkasim):

thanks again for your help

OpenStudy (anonymous):

u sub for this one

OpenStudy (anonymous):

\(u=5^x, dy =\ln(5)5^x\) should take care of it

zepdrix (zepdrix):

This is what he meant to type right there ^ \[\Large\bf\sf du=\ln(5)5^x\;dx\] You remember how to find the derivative of an exponential function like 5^x?

OpenStudy (mkmkasim):

Yeah that's what I forgot how to do

zepdrix (zepdrix):

Ooo did they fix the goodies? Equation tool and drawing tool are working??